A car of mass 100 kg moves up an inclination of 30 ⁰ at a constant speed of 20m/s calculated the power developed by an engin if the cofficient of friction is 0.2 value of acceleration due to gravity is 9.8
 

Grade:11

2 Answers

Avishek Bhagat
3 years ago
Dear Student,
answer to your problem is given below:-

As, Power=F.V
And v = 20m/s
F is force given by engine to overcome forces trying to stop car
As v is constant so net force by engine is equal to force trying to stop it
Namely Gravity and friction
F = mgsinx+ friction = 100*9.8*sin30⁰+0.2*100*9.8
F=686 N
Power = 686*20 = 13720 watts

askIITians Expert,
Yuvraj Singh
Vibekjyoti Sahoo
2 years ago
As, Power=F.V
And v = 20m/s
F is force given by engine to overcome forces trying to stop car
As v is constant so net force by engine is equal to force trying to stop it
Namely Gravity and friction
F = mgsinx+ friction = 100*9.8*sin30⁰+0.2*100*9.8
F=686 N
Power = 686*20 = 13720 watts